# Brownian dynamics simulations of electric double-layer capacitors with tunable metallicity

### v1

@@ I Introduction
The interface between metals and liquid electrolytes plays a key role in electrochemical energy storage devices (batteries, electric double layer capacitors), electrocatalysis, or electrochemical sensing. In these contexts, the interfacial properties result from the mutual influence of the electron density inside the metal and of the charge distribution due to the solvent molecules and the ions inside the liquid. This interplay controls in particular the equilibrium properties such as the ionic densities and electrostatic potential profiles, or the capacitance that links the charge accumulated on both sides of the interface to the potential drop across it. From the theoretical point of view, the predictions of equilibrium properties generally rely on the mean-field Poisson-Boltzmann theory, its linearized limit (Debye-Hückel theory), or its extensions parsons_electrical_1990; kornyshev_double-layer_2007; goodwin_mean-field_2017; baskin_improving_2017. More advanced liquid state theories such as integral equations, field theory or classical Density Functional Theory (cDFT) have been developed to include in particular electrostatic correlations kjellander_correlation_1984; zhou_image_2024, the finite size of the ions (e.g. within the primitive model of electrolytes consisting of charged hard spheres) cats_primitive_2021; cats_in-plane_2023, the explicit polarization of the solvent in addition to the ionic densities levy_dielectric_2012; levy_dipolar_2013; bruch_incorporating_2024; bruch_variational_2025 or even molecular features of the solvent around the ions jeanmairet_molecular_2013; jeanmairet_study_2019. For dynamical properties, most results have been obtained at the same level of description as Poisson-Boltzmann theory (point ions in an implicit solvent interacting via the mean-field electrostatic potential), within Poisson-Nernst-Planck theory. This model allowed in particular to investigate the charging dynamics in capacitors bazant_diffuse_2004; janssen_transient_2018; palaia_charging_2025; palaia_poisson-nernst-planck_2025 or their frequency-dependent impedance barbero_theory_2008; antonova_ambipolar_2020. Here again, progress e.g. with time-dependent cDFT allowed to analyze the effect of some physical ingredients such as the finite size of the ions or electrokinetic couplings on the dynamical response of capacitors kilic_steric_2007; jiang_time-dependent_2014; asta_lattice_2019; ma_dynamic_2022.
@@
In the last decades, these theoretical predictions were complemented by simulations at various levels of description. Ab initio or hybrid QM/MM molecular dynamics (MD) simulations provide the most accurate description but entail a large computational cost that limits the system size and trajectory length that can be simulated bonnet_first-principle_2012; elliott_qmmd_2020; le_modeling_2021; takahashi_accelerated_2022; gross_ab_2022; sakong_structure_2022; gross_challenges_2023; grisafi_predicting_2023. Classical all-atom and coarse-grained MD simulations offer a good trade-off between accuracy of the description and computational cost, and the implementation in generic or dedicated open source simulation packages weik_espresso_2019; nguyen_incorporating_2019; marin-lafleche_metalwalls_2020; coretti_metalwalls_2022; ahrens-iwers_electrode_2022 of methods to sample systems in a statistical ensemble where the voltage between the electrodes is imposed siepmann_influence_1995; reed_electrochemical_2007 have enabled their use for a variety of electrode/electrolyte interfaces. We refer the reader e.g. to Refs. scalfi_microscopic_2021 and jeanmairet_microscopic_2022 for recent reviews on these topics. Such simulations allow in particular to compute the capacitance or the frequency-dependent impedance of electrochemical cells and to correlate these observables with the microscopic properties of the confined liquid limmer_charge_2013; scalfi_charge_2020; pireddu_frequency-dependent_2023; pireddu_impedance_2024.
@@
In order to further decrease the computational cost of particle based simulations, implicit solvent models, where the effect of the solvent on ion-ion interactions is captured by the solvent permittivity and the ions evolve according to Langevin or Brownian dynamics (BD), instead of Newton’s equation of motion. The effects of the confining dielectric walls or perfect conductors and of voltage are then introduced as external potentials or enter in the ion-ion interactions, in order to model driven electrolytes confined in slit pores or capacitors under an applied voltage arnold_electrostatics_2002; tyagi_iterative_2010; breitsprecher_electrode_2015; nguyen_incorporating_2019; maxian_fast_2021; jimenez-angeles_surface_2023; dos_santos_modulation_2023; pogharian_electric_2024. For perfect conductors, efficient simulations at this level of description, taking into account periodic boundary conditions inherent to particle-based simulations in condensed matter systems, have been proposed within the Green’s function formalism to compute effective ion-ion interactions dos_santos_simulations_2017; girotto_simulations_2017; malossi_simulations_2020; telles_efficient_2024. An alternative approach for BD simulations, combining the description of explicit constant-potential electrodes using in MD simulations and ions with an implicit solvent described by its permittivity has also been proposed by Cats et al. cats_capacitance_2022. Dynamical properties, such as the electric double layer relaxation after a charge transfer event grun_relaxation_2004, the charge fluctuations of an electrode due to the interfacial ion dynamics and redox reaction katelhon_simulation-based_2012; krause_brownian_2014, or the frequency- and field-dependent response of confined electrolytes hoang_ngoc_minh_frequency_2023 have also been investigated using BD simulations.
@@
Even though most theoretical and simulation studies to date have been limited to perfect metals, the importance of charge screening inside the metal has long been recognized. At the fundamental level, this phenomenon arises from the kinetic energy of the electrons, which has to be described using quantum mechanics. Thomas-Fermi (TF) theory thomas_calculation_1927; fermi_metodo_1927 introduces this within a simple kinetic energy functional. Within a linearization approximation, the electrostatic potential then satisfies an equation identical to the Debye-Hückel equation for electrolytes, with the screening length replaced by the so-called Thomas-Fermi screening length, $l_{\rm TF}$, that depends on the density of states of the metal. This behavior can also be understood in terms of non-local electrostatics, via a wave vector dependent permittivity that depends on $l_{\rm TF}$. TF theory has been used to investigate the consequence of such screening inside the metal on ion-wall, ion-ion interactions or interfacial capacitance kornyshev_image_1977; vorotyntsev_electrostatic_1980; rochester_interionic_2013; lee_quantum_2016; lee_ionimage_2016; hedley_what_2025 (note that the electrostatic problem is the same as the one for an ion in a slit pore surrounded by an electrolyte treated at the Debye-Hückel level, that was solved analytically in Ref. levin_electrostatics_2006). The consequences on the collective interfacial behavior has been demonstrated experimentally, for example on the capillary freezing of ionic liquids confined depending on the metallicity of the substrate comtet_nanoscale_2017, which could be rationalized within TF theory kaiser_electrostatic_2017, or the kinetics of electron transfer at twisted graphene bilayers yu_tunable_2022. Several methods have also been proposed to introduced TF theory in atomistic simulations, to examine the consequences of screening inside the metal on the interfacial properties scalfi_charge_2020; schlaich_electronic_2022; goloviznina_accounting_2024; nair_induced_2025; nair_ions_2025; coello_escalante_microscopic_2024. They suffer however from the above-mentioned limitations in terms of system size, salt concentration and trajectory length due to the computational cost of molecular simulations.
@@ II.1 System
The total charge density can be written as $\rho^{\rm tot}(\mathbf{r})=\rho_{\rm l}(\mathbf{r})+\rho_{\rm ions}(\mathbf{r})+\rho_{\rm r}(\mathbf{r})$, even though only one of the three terms contributes in each region. It is the source of an electrostatic potential $\phi(\mathbf{r})$ satisfying the Poisson equation $\Delta\phi=-\rho^{\rm tot}/\varepsilon_{0}\varepsilon$ in all regions of space ($\varepsilon=1$ inside the electrodes and $\varepsilon_{\rm s}$ in the electrolyte), with $\Delta$ the Laplacian operator, as well as continuity of $\phi$ and $\varepsilon\phi^{\prime}$ at the interfaces between regions. The values of the potentials for $z\to\pm\infty$ will be discussed in the following. Since both ions and electrons are part of the system, the total electrostatic energy is
@@ II.2 Born-Oppenheimer description
which accounts for the reorganization of the charge distribution within the electrodes as the ions move. This is analogous to what is done in ab initio molecular dynamics, via the Hellmann-Feynman theorem, or to the case of explicit colloids with small ions treated at the DFT level, see Ref. lowen_ab_1992.
@@ II.2.2 Electrostatic potential
Following Ref. dos_santos_simulations_2017 for the case of a perfect metal ($l_{\rm TF}=0$), we solve the problem in reciprocal space in the $x$ and $y$ directions by looking for a Green’s function of the form
@@ II.2.3 Electrode charge
and we also used the electroneutrality of the electrolyte to introduce $\phi_{\Delta\Psi}$ in the second equality. Note that Eq. 28 can be interpreted as an equivalent circuit, with two capacitors each with capacitance per unit area $\varepsilon_{0}/l_{\rm TF}$ accounting for the effect of the TF metals in series with one with capacitance per unit area $\varepsilon_{0}\varepsilon_{\rm s}/L$ accounting for the effect of the dielectric slab (see also Ref. scalfi_semiclassical_2020). For $l_{\rm TF}=0$, Eq. II.2.3 reduces to the result derived for an ideal metal in Ref. girotto_simulations_2017.
The effective length $L_{\rm eff}$ defined in Eq. 21 can thus be understood as the width of an equivalent dielectric slab, and $E_{\rm eff}=-\Delta\Psi/L_{\rm eff}$ is the corresponding uniform electric field in the slab induced by the applied voltage. This effective length is also consistent with the one introduced to describe the lateral decay of the charge induced by an ion in a single TF electrode, proposed recently in Ref. nair_induced_2025 as an extension of earlier work by Vorotyntsev and Kornyshev vorotyntsev_electrostatic_1980, even though here the distance is the one between the two electrodes (hence the factor of two for their contribution) and not an ion-surface distance. Finally, Eq. II.2.3 can be rearranged as
@@
where we have introduced the dipole of the electrode charge distribution $M_{\rm elec}=QL_{\rm eff}$ and the solvent dipole $M_{\rm solv}=\varepsilon_{0}\varepsilon_{\rm s}L_{x}L_{y}LE$, with $E=-\Delta\Psi/L$. This underlines the various contributions to the screening of the electric field throughout the system, as discussed in the case of $l_{\rm TF}=0$ for explicit electrodes and solvent (without ions) pireddu_frequency-dependent_2023.
@@ II.2.5 Partition function, average charge and capacitance
where $\beta=1/k_{B}T$ with $k_{B}$ Boltzmann’s constant, and the corresponding thermodynamic potential is the free energy $\mathcal{F}[\Delta\Psi]=-k_{B}T\ln\mathcal{Z}[\Delta\Psi]$ (see Ref. scalfi_charge_2020). The average charge, for a given voltage, follows straightforwardly from Eq. II.2.3:
@@
similarly to what was found for simulations with explicit electrodes in the constant-potential ensemble scalfi_charge_2020. This shows that the free energy change during the charge of the capacitor from zero voltage to finite voltage $\Delta\Psi_{\rm max}$ is
@@
For a purely capacitive system, this quantity is independent of voltage and equal to the integral capacitance $C_{\rm int}=\left\langle Q\right\rangle/\Delta\Psi$. However this is not necessarily the case in general. Eq. 37 is a fluctuation-dissipation relation relating the response of the average electrode charge to a change in voltage to the equilibrium fluctuations of the electrode charge at fixed voltage. These fluctuations arise from two contributions: the thermal fluctuations of the ions, as obvious from the second term in the last line of Eq. 37, and the thermal fluctuations of the implicit electrons and the solvent, which are suppressed in the Born-Oppenheimer description and embedded in $C_{0}$, via the Thomas-Fermi screening length $l_{\rm TF}$ and the relative permittivity $\varepsilon_{\rm s}$, respectively (see Eq. 28). This result is analogous to the one obtained in Ref. scalfi_charge_2020 with explicit electrode atoms and electrolyte (including the solvent). In the latter case, the explicit expression for the contribution of the electrolyte-free capacitor has the same physical origin as $C_{0}$ but is not as simple as Eq. 28, as it depends on the microscopic details of the electrodes.
@@ II.3.1 Short-range interactions
which depends only on $z$, where $V_{\mathrm{w}}$ is the so-called Steele potential, obtained by integrating LJ interactions over atomic planes. The resulting potential includes short-range repulsion and an attractive well leading to adsorption steele_physical_interaction_of_gases_1973; steele_interaction_of_rare_gas_1978; magda_molecular_1985; magda_erratum_1986.
@@ II.3.2 Effective many-body potential
The electrostatic energy $U_{\rm Coul}^{\rm ions}[\{\mathbf{r}_{i}\}]$ can be split into two contributions: one due to direct Coulomb interactions between the ions without electrodes (corresponding to the limit of infinite interelectrode distance $L\to\infty$), which can be efficiently calculated under 2D periodic boundary conditions using Ewald summation techniques mazars_yukawa_2007a; mazars_yukawa_2007b, and another due to the polarization of the electron charge distribution inside the metal. Specifically, we write

### v2

@@ I Introduction
The interface between metals and liquid electrolytes plays a key role in electrochemical energy storage devices (batteries, electric double layer capacitors), electrocatalysis, or electrochemical sensing. In these contexts, the interfacial properties result from the mutual influence of the electron density inside the metal and of the charge distribution due to the solvent molecules and the ions inside the liquid. This interplay controls in particular the equilibrium properties such as the ionic densities and electrostatic potential profiles, or the capacitance that links the charge accumulated on both sides of the interface to the potential drop across it. From the theoretical point of view, the predictions of equilibrium properties generally rely on the mean-field Poisson-Boltzmann theory, its linearized limit (Debye-Hückel theory), or its extensions Parsons (1990); Kornyshev (2007); Goodwin et al. (2017); Baskin and Prendergast (2017); Mazur et al. (2024). More advanced liquid state theories such as integral equations, field theory or classical Density Functional Theory (cDFT) have been developed to include in particular electrostatic correlations Kjellander and Marcělja (1984); Zhou et al. (2024), the finite size of the ions (e.g. within the primitive model of electrolytes consisting of charged hard spheres) Cats et al. (2021); Cats and Härtel (2023), the explicit polarization of the solvent in addition to the ionic densities Levy et al. (2012, 2013); Bruch et al. (2024, 2025) or even molecular features of the solvent around the ions Jeanmairet et al. (2013, 2019). For dynamical properties, most results have been obtained at the same level of description as Poisson-Boltzmann theory (point ions in an implicit solvent interacting via the mean-field electrostatic potential), within Poisson-Nernst-Planck theory. This model allowed in particular to investigate the charging dynamics in capacitors Bazant et al. (2004); Janssen and Bier (2018); Palaia et al. (2025a, b) or their frequency-dependent impedance Barbero et al. (2008); Antonova et al. (2020). Here again, progress e.g. with time-dependent cDFT allowed to analyze the effect of some physical ingredients such as the finite size of the ions or electrokinetic couplings on the dynamical response of capacitors Kilic et al. (2007); Jiang et al. (2014); Asta et al. (2019); Ma et al. (2022).
@@
In the last decades, these theoretical predictions were complemented by simulations at various levels of description. Ab initio or hybrid QM/MM molecular dynamics (MD) simulations provide the most accurate description but entail a large computational cost that limits the system size and trajectory length that can be simulated Bonnet et al. (2012); Elliott et al. (2020); Le et al. (2021); Takahashi et al. (2022); Groß and Sakong (2022); Sakong et al. (2022); Groß (2023); Grisafi et al. (2023). Classical all-atom and coarse-grained MD simulations offer a good trade-off between accuracy of the description and computational cost, and the implementation in generic or dedicated open source simulation packages Weik et al. (2019); Nguyen et al. (2019); Marin-Laflèche et al. (2020); Coretti et al. (2022); Ahrens-Iwers et al. (2022) of methods to sample systems in a statistical ensemble where the voltage between the electrodes is imposed Siepmann and Sprik (1995); Reed et al. (2007) have enabled their use for a variety of electrode/electrolyte interfaces. We refer the reader e.g. to Refs. 85 and 43 for recent reviews on these topics. Such simulations allow in particular to compute the capacitance or the frequency-dependent impedance of electrochemical cells and to correlate these observables with the microscopic properties of the confined liquid Limmer et al. (2013); Scalfi et al. (2020b); Pireddu and Rotenberg (2023); Pireddu et al. (2024).
@@
In order to further decrease the computational cost of particle based simulations, implicit solvent models, where the effect of the solvent on ion-ion interactions is captured by the solvent permittivity and the ions evolve according to Langevin or Brownian dynamics (BD), instead of Newton’s equation of motion. The effects of the confining dielectric walls or perfect conductors and of voltage are then introduced as external potentials or enter in the ion-ion interactions, in order to model driven electrolytes confined in slit pores or capacitors under an applied voltage Arnold et al. (2002); Tyagi et al. (2010); Breitsprecher et al. (2015); Nguyen et al. (2019); Maxian et al. (2021); Jiménez-Ángeles et al. (2023); Dos Santos et al. (2023); Pogharian et al. (2024). For perfect conductors, efficient simulations at this level of description, taking into account periodic boundary conditions inherent to particle-based simulations in condensed matter systems, have been proposed within the Green’s function formalism to compute effective ion-ion interactions Dos Santos et al. (2017); Girotto et al. (2017); Malossi et al. (2020); Telles et al. (2024). An alternative approach for BD simulations, combining the description of explicit constant-potential electrodes using in MD simulations and ions with an implicit solvent described by its permittivity has also been proposed by Cats et al. Cats et al. (2022). Dynamical properties, such as the electric double layer relaxation after a charge transfer event Grün et al. (2004), the charge fluctuations of an electrode due to the interfacial ion dynamics and redox reaction Kätelhön and Wolfrum (2012); Krause et al. (2014), or the frequency- and field-dependent response of confined electrolytes Hoang Ngoc Minh et al. (2023) have also been investigated using BD simulations.
@@
Even though most theoretical and simulation studies to date have been limited to perfect metals, the importance of charge screening inside the metal has long been recognized. At the fundamental level, this phenomenon arises from the kinetic energy of the electrons, which has to be described using quantum mechanics. Thomas-Fermi (TF) theory Thomas (1927); Fermi (1927) introduces this within a simple kinetic energy functional. Within a linearization approximation, the electrostatic potential then satisfies an equation identical to the Debye-Hückel equation for electrolytes, with the screening length replaced by the so-called Thomas-Fermi screening length, $l_{\rm TF}$, that depends on the density of states of the metal. This behavior can also be understood in terms of non-local electrostatics, via a wave vector dependent permittivity that depends on $l_{\rm TF}$. TF theory has been used to investigate the consequence of such screening inside the metal on ion-wall, ion-ion interactions or interfacial capacitance Kornyshev et al. (1977); Vorotyntsev and Kornyshev (1980); Rochester et al. (2013); Lee et al. (2016); Lee and Perkin (2016); Hedley et al. (2025) (note that the electrostatic problem is the same as the one for an ion in a slit pore surrounded by an electrolyte treated at the Debye-Hückel level, that was solved analytically in Ref. 58). The consequences on the collective interfacial behavior has been demonstrated experimentally, for example on the capillary freezing of ionic liquids confined depending on the metallicity of the substrate Comtet et al. (2017), which could be rationalized within TF theory Kaiser et al. (2017), or the kinetics of electron transfer at twisted graphene bilayers Yu et al. (2022). Several methods have also been proposed to introduced TF theory in atomistic simulations, to examine the consequences of screening inside the metal on the interfacial properties Scalfi et al. (2020b); Schlaich et al. (2022); Goloviznina et al. (2024); Nair et al. (2025b, a); Coello Escalante and Limmer (2024). They suffer however from the above-mentioned limitations in terms of system size, salt concentration and trajectory length due to the computational cost of molecular simulations.
@@ II.1 System
The total charge density can be written as $\rho^{\rm tot}(\mathbf{r})=\rho_{\rm l}(\mathbf{r})+\rho_{\rm ions}(\mathbf{r})+\rho_{\rm r}(\mathbf{r})$, even though only one of the three terms contributes in each region. It is the source of an electrostatic potential $\phi(\mathbf{r})$ satisfying the Poisson equation $\Delta\phi=-\rho^{\rm tot}/\varepsilon_{0}\varepsilon$ in each region of space ($\varepsilon=1$ inside the electrodes and $\varepsilon_{\rm s}$ in the electrolyte), with $\Delta$ the Laplacian operator, as well as continuity of $\phi$ and $\varepsilon\phi^{\prime}$ at the interfaces between regions. The values of the potentials for $z\to\pm\infty$ will be discussed in the following. Since both ions and electrons are part of the system, the total electrostatic energy is
@@ II.2 Born-Oppenheimer description
which accounts for the reorganization of the charge distribution within the electrodes as the ions move. This is analogous to what is done in ab initio molecular dynamics, via the Hellmann-Feynman theorem, or to the case of explicit colloids with small ions treated at the DFT level, see Ref. 63.
@@ II.2.2 Electrostatic potential
Following Ref. 23 for the case of a perfect metal ($l_{\rm TF}=0$), we solve the problem in reciprocal space in the $x$ and $y$ directions by looking for a Green’s function of the form
@@ II.2.3 Electrode charge
and we also used the electroneutrality of the electrolyte to introduce $\phi_{\Delta\Psi}$ in the second equality. Note that Eq. 28 can be interpreted as an equivalent circuit, with two capacitors each with capacitance per unit area $\varepsilon_{0}/l_{\rm TF}$ accounting for the effect of the TF metals in series with one with capacitance per unit area $\varepsilon_{0}\varepsilon_{\rm s}/L$ accounting for the effect of the dielectric slab (see also Ref. 83). For $l_{\rm TF}=0$, Eq. II.2.3 reduces to the result derived for an ideal metal in Ref. 27.
The effective length $L_{\rm eff}$ defined in Eq. 21 can thus be understood as the width of an equivalent dielectric slab, and $E_{\rm eff}=-\Delta\Psi/L_{\rm eff}$ is the corresponding uniform electric field in the slab induced by the applied voltage. This effective length is also consistent with the one introduced to describe the lateral decay of the charge induced by an ion in a single TF electrode, proposed recently in Ref. 72 as an extension of earlier work by Vorotyntsev and Kornyshev Vorotyntsev and Kornyshev (1980), even though here the distance is the one between the two electrodes (hence the factor of two for their contribution) and not an ion-surface distance. Finally, Eq. II.2.3 can be rearranged as
@@
where we have introduced the dipole of the electrode charge distribution $M_{\rm elec}=QL_{\rm eff}$ and the solvent dipole $M_{\rm solv}=\varepsilon_{0}\varepsilon_{\rm s}L_{x}L_{y}LE$, with $E=-\Delta\Psi/L$. This underlines the various contributions to the screening of the electric field throughout the system, as discussed in the case of $l_{\rm TF}=0$ for explicit electrodes and solvent (without ions) Pireddu and Rotenberg (2023).
@@ II.2.5 Partition function, average charge and capacitance
where $\beta=1/k_{B}T$ with $k_{B}$ Boltzmann’s constant, and the corresponding thermodynamic potential is the free energy $\mathcal{F}[\Delta\Psi]=-k_{B}T\ln\mathcal{Z}[\Delta\Psi]$ (see Ref. 84). The average charge, for a given voltage, follows straightforwardly from Eq. II.2.3:
@@
similarly to what was found for simulations with explicit electrodes in the constant-potential ensemble Scalfi et al. (2020b). This shows that the free energy change during the charge of the capacitor from zero voltage to finite voltage $\Delta\Psi_{\rm max}$ is
@@
For a purely capacitive system, this quantity is independent of voltage and equal to the integral capacitance $C_{\rm int}=\left\langle Q\right\rangle/\Delta\Psi$. However this is not necessarily the case in general. Eq. 37 is a fluctuation-dissipation relation relating the response of the average electrode charge to a change in voltage to the equilibrium fluctuations of the electrode charge at fixed voltage. These fluctuations arise from two contributions: the thermal fluctuations of the ions, as obvious from the second term in the last line of Eq. 37, and the thermal fluctuations of the implicit electrons and the solvent, which are suppressed in the Born-Oppenheimer description and embedded in $C_{0}$, via the Thomas-Fermi screening length $l_{\rm TF}$ and the relative permittivity $\varepsilon_{\rm s}$, respectively (see Eq. 28). This result is analogous to the one obtained in Ref. 84 with explicit electrode atoms and electrolyte (including the solvent). In the latter case, the explicit expression for the contribution of the electrolyte-free capacitor has the same physical origin as $C_{0}$ but is not as simple as Eq. 28, as it depends on the microscopic details of the electrodes.
@@ II.3.1 Short-range interactions
which depends only on $z$, where $V_{\mathrm{w}}$ is the so-called Steele potential, obtained by integrating LJ interactions over atomic planes. The resulting potential includes short-range repulsion and an attractive well leading to adsorption Steele (1973, 1978); J. et al. (1985, 1986).
@@ II.3.2 Effective many-body potential
The electrostatic energy $U_{\rm Coul}^{\rm ions}[\{\mathbf{r}_{i}\}]$ can be split into two contributions: one due to direct Coulomb interactions between the ions without electrodes (corresponding to the limit of infinite interelectrode distance $L\to\infty$), which can be efficiently calculated under 2D periodic boundary conditions using Ewald summation techniques Mazars (2007a, b), and another due to the polarization of the electron charge distribution inside the metal. Specifically, we write